# Enhancing Automated Loop Invariant Generation for Complex Programs with Large Language Models

### v1

@@ 3.2 Prompts Design
To provide a more formal description of the process, we formulate the invariants generation task as follows: let $\mathbb{M}$ denote the LLMs, and $\mathbf{C}$ represent the code with a maximum of $n$ lines. Each line location in the code is denoted by $l_{0},l_{1},\cdots,l_{n}$, which may correspond to a loop or data structure. $p_{i}$ represents the data structure retrieved and $loop_{n}o$, $[v_{0},v_{1},\cdots,v_{j}]$ for the order number of the loop and the key variables of the loop separately. For each line $l_{i}$, $A_{l_{i}}$ denotes the result of the query specific to line $l_{i}$, while $I_{0}^{l_{i}},I_{1}^{l_{i}},\cdots,I_{m}^{l_{i}}$ represent the target invariants at location $l_{i}$ that point to the loop. Therefore, the current problem becomes: Given a program $\mathbf{C}$ with a maximum number of lines $n$, give loop invariants $I_{0}^{l_{i}},I_{1}^{l_{i}},\cdots,I_{m}^{l_{i}}$.
Drawing from various prompting techniques Liu et al. (2023); Wei et al. (2022); Yao et al. (2023b); Yang et al. (2024), we designed a structured prompt framework, as illustrated in Figure 2. Our approach primarily employs two types of prompts: generating predicates and loop invariants separately. The pre-loop lines are used exclusively when generating loop invariants, and at this stage, the invariants from previous loops must be provided. In other informational lines, prompts for generating predicates and prompts for generating invariants utilize their respective keywords and examples.
Components Contents
Description Description of the predicates/invariants generation task for specific loop
Analysis Demands of how to abstract the properties of the data structure/loop: Analyse the relationship between variables; $\cdots$
@@
Code $\mathbf{C}$
Location $l_{i}$
Information $\{loop_{no},[v_{0},v_{1},\cdots,v_{j}]\}$
[…]
Formatting Demands for the response formatting
Figure 2: Main prompts
Figure 3: An illustrative example of C code from FFmpeg
3.3 An illustrative example
We will use C language and ANSI/ISO C Specification Language(ACSL) to describe the loop invariants of the C program as the research object, which can be automatically checked by the verification tool Frama-C Kirchner et al. (2015) and its plugin WP. In ACSL, all the information needed is implemented as the annotation which begins with \\@ or is wrapped with \*@...*\. loop invariants are annotated beginning with loop invariant, and assert represents for assertion of certain properties at a particular program point.
Consider the program fragment illustrated in Fig 3 from the widely-used FFmpeg FFmpeg (2023) library, an open-source multimedia framework written in C. For this program fragment, we will focus on its main components. The VolDetectContext structure is employed to store audio sample values for plotting frequency distributions. It contains a single member, histogram[], which is an array of unsigned 64-bit integers used to represent the distribution of audio sample frequencies. The filter_frame function processes audio filters within FFmpeg. It iterates through each audio plane of the input audio frame (AVFrame) and adjusts the corresponding count in vd->histogram, thus forming a nested loop structure. The expression pcm[i] + 0x8000 shifts the sample values (typically ranging from -32768 to 32767) to a 0 to 0xFFFF range for indexing purposes.
3.4 Extraction Phase
In this phase, ACInv mainly applies static analysis techniques on program $\mathbf{C}$ and returns both data structure and loop information. ACInv obtains variable information for each layer of the loop by traversing the loop body and annotate the data structure used. Then ACInv sorts the loop by the control flow of programs. Each loop is assigned a sequence number, adhering to the following rules: for nested loops, the innermost loop is assigned the smallest number, while for parallel loops, the first encountered loop receives the smallest number. For each loop, ACInv captures key variable information, which is specifically, all assigned variables and parameters passed as pointers in function calls. Variable information can help LLM focus its attention on key parts, assisting in generating invariants. The sorting of loops is to further decompose the program, step-by-step generating invariants for the program.
For the example in Fig 3, ACInv will record the declaration of the structure VolDetectContext and its location information, and then traverse the program body. After completing the traversal of the loop body, it assigns sequence number 1 to the inner loop with line numbers 23-24 and extracts key variables: {i, vd}; then assigns sequence number 2 to the outer loop with line numbers 20-25 and extracts key variables: {plane, pcm[]}, as well as {i, vd} inherited from the inner loop.
3.5 Generation Phase
In this phase, ACInv will leverage the information extracted in the previous phase to generate loop invariants by querying LLMs. First, ACInv will embed the extracted data structure into the data structure prompts in this step and query the LLM to obtain predicate templates that may be used for this data structure. In prompts, we will emphasize that these predicates need to be able to abstract the data structure itself, and in addition, provide some other properties about this data structure. In the syntax of ACSL, we can use predicate to define predicates related to non-recursive data structures, and inductive to define predicates related to recursive data structures. In order to help LLM understand predicate generation, we provided some examples to LLM,. For non-recursively defined data structures, we provide a classic example: the stack, and linked list for the recursively defined data structures. The details of the examples can be found in appendix A.
@@ 3.5 Generation Phase
Then, combined with the possible predicates, ACInv embeds the program $\mathbf{C}$ and the loop information into the invariants prompt to query the LLM. Based on the order of the loops, ACInv will sequentially query the LLM for the loop with the smallest ordinal number, and require the LLM to generate invariants for that loop. Within each loop, we embed the processed program into prompts and require the LLM to analyze the following: (1) the range of values of the annotated variables; (2) Possible logical relationships between these variables and constants; (3) Does the data structure variable satisfy the predicate.
@@
Figure 3 illustrates the majority of the process for this step. Initially, two predicates are generated based on the data structure, which represents whether it is a valid VolDetectContext and the non-negativity of the array histogram[]. Using these predicates, ACInv first derives an invariant for Loop No. 1, which is an inner loop. Once this invariant is obtained, all invariants within this loop are forwarded to the next stage for enhancement. Following this, the enhanced invariant is used as a reference for generating the invariant for Loop No. 2 by querying the LLM.
@@ 3.6 Augmentation Phase
In this phase, ACInv focuses on refining the generated invariants by either strengthening or weakening them. Unlike the previous phase, this stage is loop level, which means that for each individual loop, the generated invariant will first be sent to this stage to try to be enhanced, and then the enhanced invariant group will be sent back to the generation phase as a reference for generating invariant in subsequent loops. It begins by utilizing an evaluator to assess the correctness of the invariants and then calls upon an optimizer to take appropriate actions based on the evaluation.
If an invariant is deemed correct, ACInv will attempt to "strengthen" it. In the context of invariants, this typically involves narrowing the range of program variable states represented, resulting in a more precise invariant. Conversely, if the invariant is found to be incorrect, ACInv seeks to "weaken" it. This process involves expanding the range of program variable states to encompass the correct set of states within the revised invariant. By broadening the scope of the incorrect invariant, ACInv aims to adjust it to better reflect the valid program states.
In practice, before entering the augmentation phase, the invariant generated by ACInv for Loop No. 2 in Figure 3 is 0 <= plane < nb_planes. However, during the execution of the loop, although the outer loop condition is plane < nb_planes, the final iteration may satisfy plane == nb_planes. Thus, this invariant does not cover all possible program states. The evaluator can identify this issue and the optimizer can adjust the invariant by expanding its range to 0 <= plane <= nb_planes, ultimately providing the refined invariant.

### v3

@@ 3.2 Problem description
To provide a more formal description of this problem, we formulate the invariants generation task as follows: let $\mathbb{M}$ denote the LLMs, and $\mathbb{P}$ represent the program with a maximum of $n$ lines. Each line location in the code is denoted by $l_{0},l_{1},\cdots,l_{n}$, which may correspond to a line in a loop or data structure.
Let $d_{num}\in\mathcal{D}$ denote the indexed data structure instance retrieved by ordinal numnum, while $loop_{num}\in\mathcal{L}$ corresponds to an independent loop structure identified with the same indexing scheme, where $\mathcal{D}$ and $\mathcal{L}$ is the set of corresponding information. For each loop $loop_{num}$, assuming its specific position in the program $\mathbb{P}$ starts at line $l_{i}$, where $i\in\{1,n\}$, we define:
  • $\mathbf{P}$ as the prompts, which mainly include four prompts, predicate prompt as $\mathbf{P}_{ds}$, invariant prompt as $\mathbf{P}_{inv}$, evaluator prompt as $\mathbf{P}_{ev}$, and optimizer prompt as $\mathbf{P}_{op}$,
  • $\mathbb{M}(\mathbf{P}(loop_{num},l_{i},\mathbb{P},\mathcal{D}))$ as the query result specific to $(loop_{num},l_{i})$, and
  • $\{I_{k}^{num}\}_{k=0}^{m_{num}}$ as the set of loop invariants at $l_{i}$.
Therefore, the current problem becomes: Given a program $\mathbb{P}$ with a maximum number of lines $n$, for each loop $loop_{num}\in\mathcal{L}$ and its corresponding line $l_{i}$ where $i\in\{1,n\}$, generate loop invariants $\{I_{k}^{num}\}_{k=0}^{m_{num}}\in\mathcal{I}$.
Figure 3: Diagram of properties
To formalize our requirements for the generated invariants’ properties, we employ Fig. 3 to illustrate key characteristics. In this representation, distinct variables correspond to different dimensions. The value ranges of these variables and their interrelationships collectively delineate all reachable program states, denoted as $PS$ in the figure. For simplicity, only two variables $x$ and $y$ are shown; practical scenarios may involve higher-dimensional relationships.
The invariant $I$ we aim to generate typically constitutes an over-approximation of $PS$. Consequently, our primary requirement is correctness: the invariant must subsume all reachable program states, i.e., $PS\subseteq I$. Under this correctness guarantee, a stronger invariant is preferred, characterized by a tighter approximation closer to the true program state. Formally, smaller variable ranges yield stronger invariants. As illustrated, $I_{1}$ is stronger than $I_{0}$, whereas $I_{2}$ is weaker.
3.3 An illustrative example
Figure 4: An illustrative example of C code from FFmpeg
Consider the example illustrated in Fig 4. This program fragment is from the widely-used FFmpeg [35] library, an open-source multimedia framework written in C that facilitates the recording, conversion, and streaming of audio, video, and other multimedia files. For this program fragment, we will focus on its main components, omitting less relevant details.
The VolDetectContext structure is employed to store audio sample values for plotting frequency distributions. It contains a single member, histogram[], which is an array of unsigned 64-bit integers used to represent the distribution of audio sample frequencies. The filter_frame function processes audio filters within FFmpeg, specifically handling audio frames and updating frequency distributions. It iterates through each audio plane of the input audio frame (AVFrame) and adjusts the corresponding count in vd->histogram, thus forming a nested loop structure. The expression pcm[i] + 0x8000 shifts the sample values (typically ranging from -32768 to 32767) to a 0 to 0xFFFF range for indexing purposes. This illustrative case forms the basis for our theoretical extension in the following section, where we derive generalized insights through rigorous analytical decomposition.
3.4 Extraction Phase
The primary objective of this phase is to extract structural and loop information from the target program $\mathbb{P}$. The input of this phase is a program or code segment $\mathbb{P}$, and output the data structure information $\mathcal{D}=\{d_{num}\}$ and loop information $\mathcal{L}=\{loop_{num}\}$. To achieve this, we implement a specialized parsing module that systematically traverses the entire program. The module first extracts data structure declarations and constructs a comprehensive data structure information list, denoted as $\mathcal{D}=\{d_{num}\}$. For each identified data structure, the module retrieves all its members and determines whether the data structure is recursive. A data structure is considered recursive if it contains a member of its own type. The recursiveness of a data structure significantly influences subsequent processing stages. At this phase, we merely flag the recursive nature of data structures, with detailed processing strategies to be elaborated in later stages.
More specifically, this stage actually achieves the following operations, such as Algorithm 1:
Input : A C program $\mathbb{P}$
Output : Set $\mathcal{D}$ of data structures, Set $\mathcal{L}$ of loop information
$\mathcal{D}\leftarrow\emptyset$ ;
  // Initialize data structure set
$\mathcal{L}\leftarrow\emptyset$ ;
  // Initialize loop information set
$loopStack\leftarrow\text{empty stack}$ ;
  // Track nested loop positions
1 foreach node $n$ in the Abstract Syntax Tree (AST) of $P$ do
2 if $n$ is a struct declaration then
    $\mathcal{D}\leftarrow\mathcal{D}\cup\{(datatype,dataname))\}$ ;
       // Record struct
3
4 else if $n$ is a loop (for/while/do-while) then
      push $n$ onto loopStackloopStack ;
       // Update loop position
5    $modifiedVars\leftarrow\emptyset$ ;
6    foreach variable $v$ modified in $n$’s body do
7       $modifiedVars\leftarrow modifiedVars\cup\{v\}$
    $loopDepth\leftarrow|loopStack|$ $\mathcal{L}\leftarrow\mathcal{L}\cup\{(modifiedVars,loopDepth)\}$ ;
       // Record loop
8      pop $n$ from loopStackloopStack
9
10return ($\mathcal{D}$, $\mathcal{L}$);
Algorithm 1 The algorithm for extraction phase
This algorithm performs static analysis of C programs through abstract syntax tree (AST) traversal, extracting structural and behavioral metadata. The procedure takes a C program $\mathbb{P}$ as input and outputs two sets: $\mathcal{D}$ for data structure declarations and $\mathcal{L}$ for loop information. Initialization establishes empty sets $\mathcal{D}$ and $\mathcal{L}$, along with an empty loop stack to track nesting depth. During AST traversal, when encountering a struct declaration, the algorithm extracts its composite data type definition $\tau$ and adds $\tau$ to $\mathcal{D}$. For loop constructs (for/while/do-while), it pushes the loop node onto the stack, identifies all modified variables $V_{\text{mod}}$ within the loop body, computes the current nesting depth as the stack size, records the tuple $(V_{\text{mod}},\text{depth})$ to $\mathcal{L}$, and finally pops the loop node from the stack. The algorithm terminates by returning the tuple $(\mathcal{D},\mathcal{L})$, providing comprehensive structural definitions and loop mutation patterns essential for downstream program analysis tasks such as parallelization or dataflow verification, with $\mathcal{O}(n)$ complexity proportional to AST node count.
Subsequently, ACInv systematically traverses the control flow graph of the program, with particular emphasis on loop structures. To facilitate structural analysis and provide clear decomposition pathways for the LLM, a hierarchical decomposition strategy is employed to break down complex loop constructs. During this process, ACInv identifies all loop instances within the program and assigns unique sequence numbers organized in a loop stack method. The numbering scheme follows these principles: for nested loops, the innermost loop receives the smallest number, while for parallel loops, priority is given to the loop appearing earlier in the program text. As illustrated in Fig 4, the innermost loop is labeled as 1, the enclosing loop as 2, and if there is a subsequent loop following this nested structure, it would be assigned number 3. This sequential numbering scheme establishes an analytical framework that guides the program analysis in subsequent phases, ensuring systematic exploration of the program’s control flow.
Variable information can help LLM focus its attention on key parts, providing assistance in generating invariants. For each loop, ACInv also captures key variable information, which is specifically, all assigned variables and parameters passed as pointers in function calls. The parser will provide all variables that have undergone value changes and determine their variable types, including whether they belong to the abstract data types defined by the previously extracted data structure. All the information, along with the loop sequence number, forms the loop structure information $loop_{num}$ and is then maintained in the list $\mathcal{L}=\{loop_{num}\}$.
For the example in Fig 4, ACInv will record the declaration of the structure VolDetectContext, provide a judgment that it is not recursive, and then traverse the program body. After completing the traversal of the loop body, it assigns sequence number 1 to the inner loop with line numbers 23-24 and extracts key variables: {i, vd}; then assigns sequence number 2 to the outer loop with line numbers 20-25 and extracts key variables: {plane, pcm[]}, as well as {i, vd} inherited from the inner loop.
3.5 Generation Phase
In this phase, ACInv will leverage the information extracted in the previous phase to generate loop invariants by querying LLMs. The generation module of ACInv accepts the input $(\mathbb{P},\mathcal{D},\mathcal{L})$, which is the target program and the information of data structure and loop parsed in the extraction module, and give the output $\{{I^{\prime}}_{k}^{num}\}_{k=0}^{m}$ which is the preliminary invariant for each loop $loop_{num}\in\mathcal{L}$.
Figure 5: The prompts used in ACInv
In real-world programs, a fundamental challenge in automated invariant generation stems from the prevalence of custom data structures that frequently appear in real-world programs. These structures often represent either specialized variations of conventional data structures or, more commonly, intricate compositions tailored to specific application requirements. To address this challenge, our approach emphasizes the systematic characterization and formal specification of the properties of these data structures as a prerequisite step prior to invariant generation. This preliminary analysis enables more accurate and meaningful invariant derivation that properly reflects the semantic complexity of real-world data organizations. First, ACInv will embed the data structure information $\mathcal{D}=\{d_{num}\}$ into the predicate prompts $\mathbf{P}_{ds}$, which is illustrated in Figure 5 and query the LLM to obtain predicate templates that may be used for this data structure. In prompts, we emphasize that these predicates need to be able to abstract the data structure itself, and in addition, provide some other properties about this data structure.
A critical observation reveals that the recursiveness of custom data structures constitutes a fundamental property that significantly influences both the formal specification of these structures in ACSL and the derivation of their associated properties. This characteristic plays a pivotal role in determining the appropriate representation and analysis of data structures throughout the verification process. In the syntax of ACSL, we can use predicate to define predicates related to non-recursive data structures, and inductive to define predicates related to recursive data structures. In order to help LLM understand predicate generation, we provided some examples to LLM, which are called shots in LLM prompting. The [[Examples]] section in the predicate prompts acts as a placeholder that will be dynamically filled with specific examples. Similarly, the [[Code]] placeholder will be replaced with the actual corresponding data structure. For non-recursively defined data structures, we provide a classic example: the stack, whose C language definition is usually as Figure 6(a).
In order to help LLM understand predicate generation, we provided some examples to LLM, which are called shots in LLM prompting. For non-recursively defined data structures, we provide a classic example: the stack, whose C language definition is usually as Figure 6(a).
1struct stack_int{
2 size_t top;
3 int data[MAX_SIZE];
4}
(a) An non-recursive data structure example: stack
1struct Node{
2 int data;
3 struct Node * next;
4};
(b) A recursive data structure example: linked list
Figure 6: Two examples of data structure.
@@ 3.5 Generation Phase
Figure 7: Two examples of data structure.
@@
Then, combined with the possible predicates, ACInv embeds the program $\mathbb{P}$ and the loop information $\mathcal{L}$ into the invariants prompt $\mathbf{P}_{inv}$ to query the LLM. Based on the order of the loops, ACInv will sequentially query the LLM for the loop with the smallest ordinal number, and require the LLM to generate invariants for that loop. Within each loop, we embed the processed program into prompts and require the LLM to analyze the following: (1) the range of values of the annotated variables; (2) Possible logical relationships between these variables and constants; (3) Does the data structure variable satisfies the predicate. The [[Example]] section provides multiple program instances accompanied by their corresponding loop invariants, with particular emphasis on detailed explanations of nested loop structures and the associated data structures utilized within these examples. This section serves to illustrate both the general patterns and specific considerations for invariant generation in complex control flow scenarios and custom data structure implementations. From the responses generated by the LLM, we systematically extract the invariants corresponding to their designated positional indices. Specifically, for a given loop identified by its numerical label numnum, we obtain its corresponding invariant $\{{I^{\prime}}_{k}^{num}\}_{k=0}^{m}$. When the traversal of all loops is completed, we obtain a comprehensive set of invariants, denoted as $\mathcal{I}^{\prime}=\{\{{I^{\prime}}_{k}^{1}\}_{k=0}^{m_{1}},\{{I^{\prime}}_{k}^{2}\}_{k=0}^{m_{2}},\cdots,\{{I^{\prime}}_{k}^{num}\}_{k=0}^{m_{num}},\cdots\}$.
@@
Figure 4 illustrates the majority of the process for this step. Initially, two predicates are generated based on the data structure, which represents whether it is a valid VolDetectContext and the non-negativity of the array histogram[]. Using these predicates, ACInv first derives an invariant for Loop 1, which is an inner loop. Once this invariant is obtained, all invariants within this loop are forwarded to the next stage for enhancement. Following this, the enhanced invariant is used as a reference for generating the invariant for Loop 2 by querying the LLM.
@@ 3.6 Augmentation Phase
A prevalent challenge is that the invariants generated by LLMs often tend to be weak—for instance, trivial solutions such as $\mathcal{I}=TRUE$ are frequently produced. To address this issue while simultaneously filtering out incorrect candidates, we introduce an augmentation phase. In this phase, ACInv focuses on refining the generated invariants by either strengthening or weakening them. Building upon the previous phase where we derived the initial set of invariants $\mathcal{I}^{\prime}$ from program $\mathbb{P}$, we now aim to perform a deeper analysis to refine and enhance the invariant collection. We seek to obtain an optimized set of invariants $\mathcal{I}$ that more accurately captures the program’s behavioral properties and provides stronger verification guarantees. Unlike the previous phase, this stage is loop level, which means that for each individual loop, the generated invariant will first be sent to this stage to try to be enhanced, and then the enhanced invariant group will be sent back to the generation phase as a reference for generating the invariant in subsequent loops. For each loop $loop_{num}$, the $i$-th generated invariant ${I^{\prime}}^{num}_{i}$ will be enhanced independently.
It begins by utilizing an evaluator to assess the correctness of the invariants and then calls upon an optimizer to take appropriate actions based on the evaluation. Both the evaluator and optimizer are LLM-based. As illustrated in Fig 5, we use elaborately designed dedicated prompts and examples to implement the LLM-based evaluator and optimizer. It is important to note that these components utilize LLMs independently, both with each other and with the LLM employed in the aforementioned generation phase. As mentioned earlier, we hope to use evaluators to ensure the correctness and optimizers to guarantee a stronger invariant. However, a very normal concern is that the LLM may make mistakes. However, relevant studies [20] have shown that the existing LLMs already have a strong enough reasoning ability and can give results with relatively reliable correctness rates to a certain extent. Through multiple independent information processing operations, we can further reduce its error probability. We expect to complete an end-to-end invariant generation work based on LLM, and thus it can be applied to scenarios that tolerate certain errors.
If an invariant is deemed correct, ACInv will attempt to "strengthen" it. In the context of invariants, this typically involves narrowing the range of program variable states represented, resulting in a more precise invariant. Conversely, if the invariant is deemed to be incorrect, ACInv seeks to "weaken" it. This process involves expanding the range of program variable states to encompass the correct set of states within the revised invariant. By broadening the scope of the incorrect invariant, ACInv aims to adjust it to better reflect the valid program states. However, our weakening process incorporates a rejection mechanism, allowing the LLM to decline the weakening attempt if it determines that the invariant cannot be properly weakened while maintaining correctness. Following this step, the results are subsequently returned to the evaluator for further validation. To control the iterative refinement process, we introduce a parameter $k$ that limits the maximum number of query cycles. When this threshold is reached, invariants that remain incorrect are permanently rejected, while those that have been successfully validated are retained in the final invariant set. At this stage, we obtain the final set of invariants, denoted as $\mathcal{I}=\{\{I_{k}^{1}\}_{k=0}^{m_{1}},\{I_{k}^{2}\}_{k=0}^{m_{2}},\cdots,\{I_{k}^{num}\}_{k=0}^{m_{num}},\cdots\}$.
In practice, before entering the augmentation phase, the invariant generated by ACInv for Loop 2 in Figure 4 is 0 <= plane < nb_planes. However, during the execution of the loop, although the outer loop condition is plane < nb_planes, the final iteration may satisfy plane == nb_planes. Thus, this invariant does not cover all possible program states. The evaluator can identify this issue and the optimizer can adjust the invariant by expanding its range to 0 <= plane <= nb_planes, ultimately providing the refined invariant.
@@ 4 Experiment
4.1 Reaserch Question
In this section, we will propose several intriguing research questions concerning ACInv and conduct experiments to address these questions:
  • RQ1: Can ACInv generate correct predicate templates for complex programs with data structure and thus generate correct loop invariants?
  • RQ2: Can ACInv generate correct loop invariants for complex programs without data structure?
  • RQ3: How useful is the augmentation phase for ACInv?
4.2 Experimental Setup
Benchmark Our dataset is mainly divided into two parts: those with data structures and those without data structures. The dataset of data structure consists of three parts: LIG-MM [23, 29], SV-COMP [30, 31], and FFmpeg [35]. We will elaborate on these three datasets in more detail later. The dataset without data structure mainly comes from the benchmark from AutoSpec [22], and we will also briefly describe this dataset.
First, we selected the benchmark from LIG-MM, which is a new benchmark specifically for programs with complex data structures and memory manipulations. LIG-MM is also divided into four parts: (1) the course programs which originate from course homework; (2) SV-COMP [37], the competition on software verification; (3) Benchmark from SLING [38], which is a separation logic loop invariant inferring tool using dynamic analysis; (4) Some real-world programs from well-known software and systems. LIG-MM consists of C code and focuses on some complex data structures, especially the single-linked list, double-linked list, tree, and hash table, which can be an appropriate reference dataset. Due to the limit on the number of tokens, we selected 227 examples with relatively few lines of code for testing.
We also picked C programs from SV-COMP 2024. LIG-MM and the benchmark of AutoSpec both select some programs from SV-COMP. But unlike them, we chose different parts. We selected twenty-six programs, all of which include at least one loop and one data structure. Among these data structures, we excluded all structures containing the single-linked list, double-linked list, tree, and hash table to avoid duplication with the selected parts in LIG-MM. Therefore, these programs also include many different types of structures.
To evaluate the effectiveness of ACInv on real-world programs, we selected several program segments from the widely used FFmpeg, the audio and video processing library. Given the complexity of function calls and structures within C programs, as well as the extensive codebase, we focused on specific portions of the program. We applied simple partitioning techniques, ensuring that ACInv could concentrate on the relevant functions for analysis, and we picked the programs with at least one loop and at least one data structure.
We also selected programs from AutoSpec, an automated tool for generating ACSL specifications for C programs. AutoSpec’s benchmark assesses the tool’s ability to synthesize specifications, including loop invariants, preconditions, and postconditions of functions. Notably, some C programs in this benchmark do not contain loops, so we specifically chose a subset that includes loops for our testing.
Benchmark from AutoSpec comprises four parts, which we introduce briefly: (1) Frama-C problems[39], a repository focused on program verification challenges for Frama-C. We selected a portion containing C programs with loops. (2) SyGuS[40], which represents the invariants track from the Syntax-Guided Synthesis competition, featuring only programs with linear loop structures. (3) OOPSLA-13[41, 42], and (4) SV-COMP[37], both of which include benchmarks that may have nested or multiple loops. All these benchmarks consist of C code, primarily focused on numerical programs without explicit structural declarations.